# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v2

@@ 4.2 Analysis of FFT Algorithm
Theorem 4.3 (The Analysis of FFT-based Approximation with Tunable Relative Error).
Fix a distinct pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}$ be a local randomizer that satisfies Assumption 2.3. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ Fix target knobs $\eta_{\mathrm{main}},\,\eta_{\mathrm{trunc}},\,\eta_{\mathrm{disc}},\,\eta_{\mathrm{alias}}\in\mathbb{R}.$ Let $\varepsilon_{n}\to 0$ be a sequence of privacy levels such that $\varepsilon_{n}=\Theta\bigl(\sqrt{\log n/n}\bigr).$ We denote by $L_{n}$ and $U_{n}$ the lower and upper bounds from Theorem 5, that is, $L_{n}:=P(-c;x_{1})-\mathrm{e}^{\varepsilon}P(c;x_{1}^{\prime})-\delta^{\mathrm{low}}_{\mathrm{err}}$ and $U_{n}:=P(c;x_{1})-\mathrm{e}^{\varepsilon}P(-c;x_{1}^{\prime})+\delta^{\mathrm{up}}_{\mathrm{err}}$ We define our numerical approximation as the midpoint $\widehat{D}_{n}:=\frac{U_{n}+L_{n}}{2}.$
@@
See Appendix A.7 for the full proof. In particular, in the regime $\varepsilon_{n}=\Theta(\sqrt{\log n/n})$, the theorem shows that we can prescribe a target relative error level $\eta>0$ and choose the numerical knobs so that the relative error becomes $O(\eta)$, which leads to an overall running time $\widetilde{O}\!\left(\frac{n}{\eta}\right),$ up to polylogarithmic factors in $n$ and $1/\eta$. We will provide details on parameter selection and implementation in Appendix C.
@@ 5 Experiments
(a) Upper ($\circ$) and lower ($\square$) shuffled mechanism’s DP $\delta$ for $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ of 3-$R$ with $\varepsilon_{0}=2$ (left) and the Gaussian mechanism with $\sigma=2$ (right), both with annotated $\varepsilon_{n}$; the dotted line indicates $\alpha/n$.
(b) Relative error as a function of the accuracy knob $\eta_{\mathrm{main}}$ with annotated runtime.
(c) Runtime as a function of the number of users $n$ with annotated relative error.
(d) Lower shuffle index as a function of the RMSE for mean estimation with annotated RMSE for variance and 33rd moment estimation.
@@
In this section, we complement our theoretical results with five experiments. The first two visualize the privacy curves from Theorems 3.3 and 3.4. The next two illustrate the accuracy-runtime trade-offs of our FFT-based accountant from Theorems 5 and 4.3. The final experiment compares generalized Gaussian mechanisms with $k$-$R$ in a distribution-estimation task.
Unless stated otherwise, all computations are carried out in double precision and we choose $(c,s,w^{\mathrm{in}},h,w^{\mathrm{out}})$ according to Theorem 5 and set $\eta_{\mathrm{trunc}},\eta_{\mathrm{disc}},\eta_{\mathrm{alias}}=10^{-3}$. For each mechanism we first identify the asymptotically worst-case neighboring pair $(x_{1},x_{1}^{\prime})$ and the corresponding shuffle indices $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ as in Theorem 3.3, and then run the FFT-based accountant at this pair.
@@ 5.1 Privacy Curves and Shuffle Index
We empirically evaluate the privacy guarantees of Theorem 3.3. For a fixed target constant $\alpha>0$ and a given local randomizer we set $\varepsilon_{n}=\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ and explore the behavior of the shuffled mechanism’s hockey-stick divergence $\delta_{n}$ as a function of $n$ for two mechanisms. That is, we use the blanket-based upper and lower bounds on the true shuffled hockey-stick divergence: the Balle et al. upper bound (Lemma 2.1) and the Su et al. lower bound (Theorem 2.2), both evaluated via our FFT accountant at the asymptotically worst-case pair.
The left panel of Figure 1(a) shows the results for the case of $3$-$R$ with $\varepsilon_{0}=2$. For each $n$ we plot the upper and lower bounds on the true hockey-stick divergence $\delta_{n}$ of the shuffled mechanism corresponding to $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$. Across the entire range of $n$ the two curves are practically indistinguishable. This behavior is exactly what Theorem 3.4 predicts for $k$-$R$ with $k\geq 3$. On the theoretical side, the lower and upper shuffle indices coincide, $\underline{\chi}_{\mathrm{lo}}=\underline{\chi}_{\mathrm{up}}$, so the asymptotic band in Theorem 3.3 collapses. Empirically, we see that this collapse already occurs for moderate values of $n$, confirming that the blanket-divergence analysis is effectively optimal for this mechanism.
The right panel of Figure 1(a) reports the same experiment for the Gaussian local randomizer with noise parameter $\sigma=2$. Here the upper and lower curves form a visible but relatively narrow band. This matches our theoretical findings: the Gaussian mechanism does not satisfy the structural condition of Theorem 3.4, so the shuffle indices $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ need not coincide. Nevertheless, their ratio stays close to one (above $0.7$ in our parameter range), and the band remains tight across all $n$.
@@ 5.2 Accuracy and Runtime of the FFT Accountant
Figure 1(b) shows the effect of the parameter $\eta_{\mathrm{main}}$ on both accuracy and runtime. For each value of $\eta_{\mathrm{main}}$ we run the FFT-based algorithm for a fixed $(n,\varepsilon)$. Theorem 5 provides rigorous lower and upper bounds on the blanket divergence, $D_{\mathrm{low}}\;\leq\;\mathcal{D}^{\mathrm{blanket}}\;\leq\;D_{\mathrm{up}},$ so the true value is guaranteed to lie in the interval $D_{\mathrm{low}},D_{\mathrm{up}}]$. As our notion of relative error, we plot the certified relative uncertainty bandwidth $\frac{D_{\mathrm{up}}-D_{\mathrm{low}}}{D_{\mathrm{up}}}.$ As predicted by Theorem 4.3, the relative error shrinks approximately linearly as $\eta_{\mathrm{main}}$ decreases, while the runtime grows roughly like $1/\eta_{\mathrm{main}}$.
In Figure 1(c) we fix $\eta_{\mathrm{main}}$ and vary the number of users $n$. For each $n$ we again compute the interval $D_{\mathrm{low}},D_{\mathrm{up}}]$ and plot both the runtime and the relative error. The relative error remains essentially flat as $n$ increases, confirming that our parameter selection keeps the relative error under control uniformly in $n$. At the same time, the runtime grows almost linearly in $n$, with only mild logarithmic deviations. This matches the theoretical $\widetilde{O}(n/\eta)$ complexity of Theorem 4.3.
@@ 5.3 Distribution Estimation
Finally, we compare the generalized Gaussian mechanisms with $k$-$R$ in a distribution estimation task. We fix the data-generating distribution to be $X\sim\mathrm{Unif}[0,1]$ and set the number of users to $n=10^{5}$. For each local mechanism, we evaluate the empirical root-mean-squared error (RMSE) of the mean estimation. Figure 1(d) plots the RMSE on the horizontal axis and the lower shuffle index $\underline{\chi}_{\mathrm{lo}}$ on the vertical axis. For generalized Gaussian mechanisms we additionally annotate each point with the RMSEs for variance and third-moment estimation, whereas for $k$-$R$ we only plot the mean-accuracy vs. shuffle-index trade-off.
@@
In this experiment we focus on $k$-$R$ with $k=3$, which we found to yield the best shuffle index among the values of $k$ under the same mean-estimation accuracy. Under this choice, the figure shows that, for the same level of mean RMSE, $3$-$R$ achieves the largest lower shuffle index in our parameter range, and hence provides the strongest privacy. However, a crucial caveat is that $3$-$R$ (and, more generally, $k$-$R$) does not natively support estimation of higher-order moments: estimating the variance or third central moment typically requires designing a separate mechanism (and hence additional privacy budget). In contrast, generalized Gaussian mechanisms simultaneously support estimation of multiple moments from the same shuffled output. Figure 1(d) reveals that the generalized Gaussian family exhibits a nontrivial dependence on the shape parameter $\beta$: the value of $\beta$ that optimizes the shuffle index varies with the target mean-estimation accuracy. In particular, in more noisy regimes (larger local noise), the Gaussian mechanism ($\beta=2$) attains the largest lower shuffle index among the generalized Gaussian family. Moreover, over a wide range of mean-accuracy levels, the Gaussian mechanism also achieves the smallest RMSE for variance and third-moment estimation among the generalized Gaussian mechanisms, as indicated by the annotations. Taken together, these results suggest that generalized Gaussian mechanisms ($\beta>1$) provide a better utility-privacy trade-off than pure local DP mechanisms. Importantly, this kind of optimization is made possible by our unified framework beyond pure local DP.
6 Limitations and Conclusion
Our analysis has a number of limitations that we highlight as directions for future work.
First, when studying the generalized Gaussian mechanism we restricted attention to the one-dimensional case. This restriction is not conceptually essential. Our asymptotic and FFT-based arguments extend in principle to higher-dimensional local randomizers; however, in high dimensions the blanket mass $\gamma$ of the Gaussian mechanism decays exponentially in the dimension. As a consequence, the lower shuffle index $\underline{\chi}_{\mathrm{lo}}$ shrinks rapidly with the dimension, and the blanket-divergence upper bound becomes essentially vacuous for practically relevant values of $n$. This is particularly problematic when analyzing locally private federated learning or shuffled DP-SGD through our framework. At the same time, the upper shuffle index $\underline{\chi}_{\mathrm{up}}$ exhibits essentially the same behavior as in the one-dimensional case, so that the gap between upper and lower shuffle indices widens with the dimension. Closing this gap in high dimensions appears to require techniques that go beyond the blanket divergence, which remains an open problem.
@@ 6 Limitations and Conclusion
Second, our moderate-deviation guarantees are formulated in terms of an asymptotically worst-case neighboring pair, that is, a pair $(x_{1}^{\ast},x_{1}^{\prime\ast})$ that attains the worst-case shuffle index in the limit $n\to\infty$. While this is natural from the point of view of our CLT-based analysis, it does not rule out the possibility that, for finite $n$, special pairs might induce slightly worse $(\varepsilon,\delta)$ trade-offs than those induced by the asymptotic worst-case pair. Establishing conditions under which the asymptotically worst-case pair is also worst-case (or near worst-case) at finite $n$ is a practically relevant problem that we leave for future work.
@@
Despite these limitations, our work is, to the best of our knowledge, the first to go beyond the pure local DP viewpoint and organize privacy amplification by shuffling via the shuffle index. Our framework provides a unified way to reason about shuffle amplification and to obtain tight $(\varepsilon,\delta)$ guarantees. We believe that these conceptual and technical advantages are valuable both for the theoretical understanding of shuffle DP and for its practical deployment in real systems.
@@ Appendix A Missing Proofs
A.1 Proofs of Lemma 3.1
See 3.1
@@
A.2 The proof of Lemma 3.2
See 3.2
We analyze the blanket divergence by deriving an asymptotic expansion based on central limit theorems (CLT). The general CLT-based results used in our analysis are stated below in a form tailored to our setting, with the level of abstraction kept to the minimum necessary. In particular, we treat the continuous case (Cont) and the bounded case (Bound) separately, using a triangular-array Edgeworth expansion in the former and a triangular-array moderate deviation theorem in the latter.
Theorem A.1 (Triangular-array Edgeworth expansion after Angst–Poly Angst and Poly (2017)).
@@
Theorem A.1 is obtained by specializing the general triangular-array Edgeworth expansion of Angst and Poly (Angst and Poly, 2017, Thm. 4.3) to the one-dimensional row-wise i.i.d. case. Their theorem covers much more general triangular arrays (including non-identically distributed and vector-valued arrays); in this paper we only need the resulting triangular-array extension of the classical Cramér-type Edgeworth expansion of Petrov (Petrov, 1975, Ch. VI, Thms. 3–4). Compared with the hypotheses of (Angst and Poly, 2017, Thm. 4.3), we adopt stronger assumptions in the spirit of Petrov: we require the existence of moments of all integer orders and replace the weak mean Cramér condition by the classical Cramér condition for the common distribution of $X_{1,n}$. When the distribution of $X_{1,n}$ does not depend on $n$, Theorem A.1 reduces to Petrov’s one-dimensional Edgeworth expansion for sums of i.i.d. random variables.
Theorem A.3 (Triangular-array moderate deviations after Slastnikov Slastnikov (1979)).
@@
Theorem A.3 is a specialization of Slastnikov’s moderate deviation theorem for triangular arrays (Slastnikov, 1979, Thm. 1.1) to the row-wise i.i.d. case, under the uniform boundedness assumption (iii). In Slastnikov’s notation, assumption (iii) implies his moment condition for any fixed $c>0$, and (i)–(iii) together yield $B_{n}^{2}=\mathrm{Var}(S_{n})=\Theta(n)$. Hence his conclusion $1-F_{n}(x)\sim 1-\Phi(x)$ holds uniformly in the moderate deviation region $0\leq x\leq c\sqrt{\log B_{n}}$ for any fixed $c>0$. Since $\log B_{n}=\tfrac{1}{2}\log n+O(1)$, this is equivalent to uniformity on $0\leq x\leq c\sqrt{\log n}$ as stated in Theorem A.3.
@@
Second, for the truncated variables we analyze the expectation appearing in the transformation formula of Lemma 3.1. In the continuous case (Cont) we apply a triangular-array Edgeworth expansion (Theorem A.1) to the normalized sums and derive a precise asymptotic expression for the expectation; this is summarized in Lemma A.6. In the discrete bounded case (Bound) we instead rely on Slastnikov’s triangular-array moderate deviation theorem (Theorem A.3); see Corollary A.7.
@@
We use the integral representation (i.e., (ii) of Lemma 3.1) [^9] We deliberately work with the integral representation (ii) in Lemma 3.1 rather than the linear form (i). In the discrete bounded case, starting from (i) would lead to terms of the form $\mathbb{E}[Z_{1}(1-F_{n-1}(u))\rho_{n}(u)]$, and controlling the contribution of the Slastnikov remainder $\rho_{n}(u)$ relative to the main term would require an explicit convergence rate for the moderate deviation error, which is not available in the version we use. By contrast, the integral representation keeps the factor $(1+\rho_{n}(x))$ inside the same tail integral and allows us to obtain a relative $o(1)$ error without an explicit rate. Conversely, in the continuous case we avoid representation (ii), because our Edgeworth-based proof works under moment assumptions only and does not assume the existence of exponential moments, which are convenient for the Chernoff-type tail bounds needed in the discrete bounded analysis. together with Bernoulli selectors. For each $n$ let $Y_{1},\dots,Y_{n}$ be i.i.d. with distribution $\mathcal{R}_{\mathrm{ref}}$, let $B_{1},\dots,B_{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ independent of $(Y_{i})_{i\geq 1}$. That is,
@@
Assume that the infima in the definitions of $\underline{\chi}_{\mathrm{lo}}$ and $\underline{\chi}_{\mathrm{up}}$ are attained by some distinct pairs $(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})$ and $(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}})$, respectively.
Step 1: $(i)\Rightarrow(ii)$. Suppose first that $\underline{\chi}_{\mathrm{up}}=\underline{\chi}_{\mathrm{lo}}$. By definition,
$\underline{\chi}_{\mathrm{lo}}=\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}}),\qquad\underline{\chi}_{\mathrm{up}}=\chi_{\mathrm{up}}(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}}).$

### v3

@@ 4.2 Analysis of FFT Algorithm
Theorem 4.3.
Fix a distinct pair $x_{1},x_{1}^{\prime}\in\mathcal{X}$ and let $\mathcal{R}\in\mathfrak{R}$. Let the reference distribution be $\mathcal{R}_{\mathrm{ref}}\in\bigl\{\mathcal{R}_{\mathrm{BG}}\bigr\}\,\cup\,\bigl\{\mathcal{R}_{x}:x\in\mathcal{X}\bigr\}.$ Fix target knobs $\eta_{\mathrm{main}},\,\eta_{\mathrm{trunc}},\,\eta_{\mathrm{disc}},\,\eta_{\mathrm{alias}}\in\mathbb{R}.$ Let $\varepsilon_{n}\to 0$ be a sequence of privacy levels such that $\varepsilon_{n}=\Theta\bigl(\sqrt{\log n/n}\bigr).$ We denote by $L_{n}$ and $U_{n}$ the lower and upper bounds from Theorem 5, that is, $L_{n}:=P(-c;x_{1})-\mathrm{e}^{\varepsilon}P(c;x_{1}^{\prime})-\delta^{\mathrm{low}}_{\mathrm{err}}$ and $U_{n}:=P(c;x_{1})-\mathrm{e}^{\varepsilon}P(-c;x_{1}^{\prime})+\delta^{\mathrm{up}}_{\mathrm{err}}$ We define our numerical approximation as the midpoint $\widehat{D}_{n}:=\frac{U_{n}+L_{n}}{2}.$
@@
See Appendix A.6 for the full proof. In particular, in the regime $\varepsilon_{n}=\Theta(\sqrt{\log n/n})$, the theorem shows that we can prescribe a target relative error level $\eta>0$ and choose the numerical knobs so that the relative error becomes $O(\eta)$, which leads to an overall running time $\widetilde{O}\!\left(\frac{n}{\eta}\right),$ up to polylogarithmic factors in $n$ and $1/\eta$. We provide details on parameter selection and implementation in Appendix D.
We briefly discuss regimes in which the algorithm can become less effective in practice. First, when the shuffle index $\chi$ is small, the relevant distributions for the privacy-amplification sum are effectively more spread out, so representing them on a discrete grid requires a larger window and more bins; heuristically, the required grid size (and hence runtime) can grow like $O(1/\chi^{2})$, making computations expensive for small $\chi$. In practice, as discussed in Section 3.4, privacy amplification by shuffling is meaningful only when the shuffle index is at least moderately large; the parameter regime of primary interest typically corresponds to larger $\chi$, where the FFT computation remains well behaved. Second, our parameter-tuning rules are asymptotic, so for small $n$ they may be miscalibrated and yield insufficient accuracy. That said, the method produces certified upper and lower bounds, so one can directly inspect the resulting gap to assess tightness and, if needed, retune the numerical parameters to achieve the desired accuracy. Finally, while floating-point round-off is not captured by the real-arithmetic analysis, prior work shows that the FFT-based computations can remain accurate down to about $10^{-10}$ when using long double precision Gopi et al. (2021), which is sufficient for the typical range of $\delta$ values of interest in DP.
@@ 5 Experiments
(a) Upper ($\circ$) and lower ($\square$) shuffled mechanism’s DP $\delta$ for $\varepsilon_{n}(\alpha,\underline{\chi}_{\mathrm{lo}})$ of 3-$R$ with $\varepsilon_{0}=2$ (left) and the Gaussian mechanism with $\sigma_{0}=2$ (right), both with annotated $\varepsilon_{n}$; the dotted line indicates $\alpha/n$.
(b) Comparison of the shuffled $\varepsilon$ as a function of $n$ between our analysis and that of Feldman et al. (2022). We use the Gaussian mechanism with $\sigma_{0}=2$ and fix $\delta=10^{-5}$.
(c) Relative error as a function of the accuracy knob $\eta_{\mathrm{main}}$ with annotated runtime.
(d) Runtime as a function of the number of users $n$ with annotated relative error.
(e) $\chi_{\mathrm{lo}}$ as a function of the RMSE for mean estimation with annotated RMSE for variance and 33rd moment estimation.
@@
In this section, we complement our theoretical results with five experiments. The first two visualize the privacy curves from Theorems 3.4 and 3.7. The next two illustrate the accuracy-runtime trade-offs of our FFT-based accountant from Theorems 5 and 4.3. The final experiment compares generalized Gaussian mechanisms with $k$-$R$ in a distribution-estimation task.
Unless stated otherwise, all computations are carried out in double precision and we choose $(c,s,w^{\mathrm{in}},h,w^{\mathrm{out}})$ according to Theorem 5 and set $\eta_{\mathrm{trunc}},\eta_{\mathrm{disc}},\eta_{\mathrm{alias}}=10^{-3}$. For each mechanism we first identify the asymptotically worst-case neighboring pair $(x_{1},x_{1}^{\prime})$ and the corresponding shuffle indices $\chi_{\mathrm{lo}}$ and $\chi_{\mathrm{up}}$ as in Theorem 3.4, and then run the FFT-based accountant at this pair.
@@ 5.1 Privacy Curves and Shuffle Index
We empirically evaluate the privacy guarantees of Theorem 3.4. For a fixed target constant $\alpha>0$ and a given local randomizer we set $\varepsilon_{n}=\varepsilon_{n}(\alpha,\chi_{\mathrm{lo}})$ and explore the behavior of the shuffled mechanism’s hockey-stick divergence $\delta_{n}$ as a function of $n$ for two mechanisms. That is, we use the blanket-based upper and lower bounds on the true shuffled hockey-stick divergence: the Balle et al. upper bound (Lemma 2.1) and the Su et al. lower bound (Theorem 2.2), both evaluated via our FFT accountant at the asymptotically worst-case pair.
The left panel of Fig. 4(a) shows the results for the case of $3$-$R$ with $\varepsilon_{0}=2$. For each $n$ we plot the upper and lower bounds on the true hockey-stick divergence $\delta_{n}$ of the shuffled mechanism corresponding to $\varepsilon_{n}(\alpha,\chi_{\mathrm{lo}})$. Across the entire range of $n$ the two curves are practically indistinguishable. This behavior is exactly what Theorem 3.7 predicts for $k$-$R$ with $k\geq 3$. On the theoretical side, the lower and upper shuffle indices coincide, $\chi_{\mathrm{lo}}=\chi_{\mathrm{up}}$, so the asymptotic band in Theorem 3.4 collapses. Empirically, we see that this collapse already occurs for moderate values of $n$, confirming that the blanket-divergence analysis is effectively optimal for this mechanism.
The right panel of Fig. 4(a) reports the same experiment for the Gaussian local randomizer with noise parameter $\sigma=2$. Here the upper and lower curves form a visible but relatively narrow band. This matches our theoretical findings: the Gaussian mechanism does not satisfy the structural condition of Theorem 3.7, so the shuffle indices $\chi_{\mathrm{lo}}$ and $\chi_{\mathrm{up}}$ need not coincide. Nevertheless, their ratio stays close to one (above $0.7$ in our parameter range), and the band remains tight across all $n$.
Next, we compare our direct analysis of the blanket divergence with the mechanism-agnostic shuffle analysis of Feldman et al. Feldman et al. (2022) of the calibrated Gaussian local randomizer Balle and Wang (2018). Fig. 4(b) shows the results. Because of the mechanism-agnostic nature and the reliance on a coarse reduction to a pure LDP surrogate, the resulting bounds incur a large constant-factor loss. Empirically, we observe that even at $n=10^{3}$, the method predicts essentially no privacy amplification, and for larger $n$ the predicted shuffled $\varepsilon$ remains $10$ times larger than our direct analysis.
@@ 5.2 Accuracy and Runtime of the FFT Accountant
Fig. 4(c) shows the effect of the parameter $\eta_{\mathrm{main}}$ on both accuracy and runtime. For each value of $\eta_{\mathrm{main}}$ we run the FFT-based algorithm for a fixed $(n,\varepsilon)$. Theorem 5 provides rigorous lower and upper bounds on the blanket divergence, $D_{\mathrm{low}}\;\leq\;\mathcal{D}^{\mathrm{blanket}}\;\leq\;D_{\mathrm{up}},$ so the true value is guaranteed to lie in the interval $D_{\mathrm{low}},D_{\mathrm{up}}]$. As our notion of relative error, we plot the certified relative uncertainty bandwidth $\frac{D_{\mathrm{up}}-D_{\mathrm{low}}}{D_{\mathrm{up}}}.$ As predicted by Theorem 4.3, the relative error shrinks approximately linearly as $\eta_{\mathrm{main}}$ decreases, while the runtime grows roughly like $1/\eta_{\mathrm{main}}$.
In Fig. 4(d) we fix $\eta_{\mathrm{main}}$ and vary the number of users $n$. For each $n$ we again compute the interval $D_{\mathrm{low}},D_{\mathrm{up}}]$ and plot both the runtime and the relative error. The relative error remains essentially flat as $n$ increases, confirming that our parameter selection keeps the relative error under control uniformly in $n$. At the same time, the runtime grows almost linearly in $n$, with only mild logarithmic deviations. This matches the theoretical $\widetilde{O}(n/\eta)$ complexity of Theorem 4.3.
@@ 5.3 Distribution Estimation
Finally, we compare the generalized Gaussian mechanisms with $k$-$R$ in a distribution estimation task. We fix the data-generating distribution to be $X\sim\mathrm{Unif}[0,1]$ and set the number of users to $n=10^{5}$. For each local mechanism, we evaluate the empirical root-mean-squared error (RMSE) of the mean estimation. Fig. 4(e) plots the RMSE on the horizontal axis and the lower shuffle index $\chi_{\mathrm{lo}}$ on the vertical axis. For generalized Gaussian mechanisms we additionally annotate each point with the RMSEs for variance and third-moment estimation, whereas for $k$-$R$ we only plot the mean-accuracy vs. shuffle-index trade-off.
@@
In this experiment we focus on $k$-$R$ with $k=3$, which we found to yield the best shuffle index among the values of $k$ under the same mean-estimation accuracy. Under this choice, the figure shows that, for the same level of mean RMSE, $3$-$R$ achieves the largest lower shuffle index in our parameter range, and hence provides the strongest privacy. However, a crucial caveat is that $3$-$R$ (and, more generally, $k$-$R$) does not natively support estimation of higher-order moments: estimating the variance or third central moment typically requires designing a separate mechanism (and hence additional privacy budget). In contrast, generalized Gaussian mechanisms simultaneously support estimation of multiple moments from the same shuffled output. Fig. 4(e) reveals that the generalized Gaussian family exhibits a nontrivial dependence on the shape parameter $\beta$: the value of $\beta$ that optimizes the shuffle index varies with the target mean-estimation accuracy. In particular, in more noisy regimes (larger local noise), the Gaussian mechanism ($\beta=2$) attains the largest lower shuffle index among the generalized Gaussian family. Moreover, over a wide range of mean-accuracy levels, the Gaussian mechanism also achieves the smallest RMSE for variance and third-moment estimation among the generalized Gaussian mechanisms, as indicated by the annotations. Taken together, these results suggest that generalized Gaussian mechanisms ($\beta>1$) provide a better privacy-utility trade-off than pure local DP mechanisms. Importantly, this kind of optimization is made possible by our unified framework beyond pure local DP.
6 Limitation and Conclusion
@@
Our moderate-deviation guarantees are formulated in terms of an asymptotically worst-case neighboring pair, that is, a pair $(x_{1}^{\ast},x_{1}^{\prime\ast})$ that attains the worst-case shuffle index in the limit $n\to\infty$. While this is natural from the point of view of our CLT-based analysis, it does not rule out the possibility that, for finite $n$, special pairs might induce slightly worse $(\varepsilon,\delta)$ trade-offs than those induced by the asymptotic worst-case pair. Establishing conditions under which the asymptotically worst-case pair is also worst-case at finite $n$ is a problem that we leave for future work.
@@
Our work is, to the best of our knowledge, the first to go beyond the pure local DP viewpoint and organize privacy amplification by shuffling via the shuffle index. Our framework provides a unified way to reason about shuffle amplification and to obtain tight $(\varepsilon,\delta)$ guarantees. We believe that these conceptual and technical advantages are valuable both for the theoretical understanding of shuffle DP and for its practical deployment in real systems.
@@ Appendix A Missing Proofs
A.1 The proof of Lemma 3.2
See 3.2
To prove Lemma 3.2, we start by deriving two alternative representations of the blanket divergence that are more amenable to asymptotic analysis.
Remark A.1 (Transformations for the blanket divergence).
We retain both transformations because the conditions required to identify the leading term in Lemma 3.2 differ between the continuous (Cont) and bounded (Bound) assumptions of the local randomizer (see Assumption 2.3). In the continuous case, we apply Theorem A.2 to the representation (3). In the bounded case (Bound), we instead apply Theorem A.4 to the representation (4). The resulting leading terms coincide, and we obtain the result.
@@
We now begin the proof of Lemma 3.2. We analyze the blanket divergence by deriving an asymptotic expansion based on central limit theorems (CLT). The general CLT-based results used in our analysis are stated below in a form tailored to our setting, with the level of abstraction kept to the minimum necessary. In particular, we treat the continuous case (Cont) and the bounded case (Bound) separately, using a triangular-array Edgeworth expansion in the former and a triangular-array moderate deviation theorem in the latter.
Theorem A.2 (Triangular-array Edgeworth expansion after Angst–Poly [1]).
@@
Theorem A.2 is obtained by specializing the general triangular-array Edgeworth expansion of Angst and Poly [1, Thm. 4.3] to the one-dimensional row-wise i.i.d. case. Their theorem covers much more general triangular arrays (including non-identically distributed and vector-valued arrays); in this paper we only need the resulting triangular-array extension of the classical Cramér-type Edgeworth expansion of Petrov [21, Ch. VI, Thms. 3–4]. Compared with the hypotheses of [1, Thm. 4.3], we adopt stronger assumptions in the spirit of Petrov: we require the existence of moments of all integer orders and replace the weak mean Cramér condition by the classical Cramér condition for the common distribution of $X_{1,n}$. When the distribution of $X_{1,n}$ does not depend on $n$, Theorem A.2 reduces to Petrov’s one-dimensional Edgeworth expansion for sums of i.i.d. random variables.
Theorem A.4 (Triangular-array moderate deviations after Slastnikov [22]).
@@
Theorem A.4 is a specialization of Slastnikov’s moderate deviation theorem for triangular arrays [22, Thm. 1.1] to the row-wise i.i.d. case, under the uniform boundedness assumption (iii). In Slastnikov’s notation, assumption (iii) implies his moment condition for any fixed $c>0$, and (i)–(iii) together yield $B_{n}^{2}=\mathrm{Var}(S_{n})=\Theta(n)$. Hence his conclusion $1-F_{n}(x)\sim 1-\Phi(x)$ holds uniformly in the moderate deviation region $0\leq x\leq c\sqrt{\log B_{n}}$ for any fixed $c>0$. Since $\log B_{n}=\tfrac{1}{2}\log n+O(1)$, this is equivalent to uniformity on $0\leq x\leq c\sqrt{\log n}$ as stated in Theorem A.4.
@@
Second, for the truncated variables we analyze the expectation appearing in the transformation formula of Lemma A.1. In the continuous case (Cont) we apply a triangular-array Edgeworth expansion (Theorem A.2) to the normalized sums and derive a precise asymptotic expression for the expectation; this is summarized in Lemma A.7. In the discrete bounded case (Bound) we instead rely on Slastnikov’s triangular-array moderate deviation theorem (Theorem A.4); see Corollary A.8.
@@
We use the integral representation (i.e., (ii) of Lemma A.1) [^9] We deliberately work with the integral representation (ii) in Lemma A.1 rather than the linear form (i). In the discrete bounded case, starting from (i) would lead to terms of the form $\mathbb{E}[Z_{1}(1-F_{n-1}(u))\rho_{n}(u)]$, and controlling the contribution of the Slastnikov remainder $\rho_{n}(u)$ relative to the main term would require an explicit convergence rate for the moderate deviation error, which is not available in the version we use. By contrast, the integral representation keeps the factor $(1+\rho_{n}(x))$ inside the same tail integral and allows us to obtain a relative $o(1)$ error without an explicit rate. Conversely, in the continuous case we avoid representation (ii), because our Edgeworth-based proof works under moment assumptions only and does not assume the existence of exponential moments, which are convenient for the Chernoff-type tail bounds needed in the discrete bounded analysis. together with Bernoulli selectors. For each $n$ let $Y_{1},\dots,Y_{n}$ be i.i.d. with distribution $\mathcal{R}_{\mathrm{ref}}$, let $B_{1},\dots,B_{n}$ be i.i.d. $\mathrm{Bernoulli}(\gamma)$ independent of $(Y_{i})_{i\geq 1}$. That is,
@@
Assume that the infima in the definitions of $\chi_{\mathrm{lo}}$ and $\chi_{\mathrm{up}}$ are attained by some distinct pairs $(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}})$ and $(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}})$, respectively.
Step 1: $(i)\Rightarrow(ii)$. Suppose first that $\chi_{\mathrm{up}}=\chi_{\mathrm{lo}}$. By definition,
$\chi_{\mathrm{lo}}=\chi_{\mathrm{lo}}(x_{1}^{\mathrm{lo}},x_{1}^{\prime\,\mathrm{lo}}),\qquad\chi_{\mathrm{up}}=\chi_{\mathrm{up}}(x_{1}^{\mathrm{up}},x_{1}^{\prime\,\mathrm{up}}).$